# KCL-PH-TH/2025-33

### v1

@@ I Introduction
The standard cosmological narrative begins with a period of slow-roll inflation [1, 2, 3], during which the dynamics are governed by a scalar field—the inflaton—rolling down a nearly flat potential. This phase provides solutions to the horizon and flatness problems, while also stretching quantum fluctuations of the inflaton field, which later seeds the formation of large-scale structures.After inflation ends,it is usually assumed that the inflationary sector reheats the universe by transferring most of its energy to Standard Model particles. This marks the beginning of the Hot Big Bang phase, initially characterized by a radiation-dominated plasma with an equation-of-state parameter $w=1/3$, which eventually transitions to matter domination with $w=0$, as the universe cools. Interestingly, there are no strong constraints on the equation of state of the universe during the period between the end of inflation and the onset of reheating. This opens up the possibility for a variety of cosmological scenarios. One such possibility is that the post-inflationary–but pre-reheating–epoch is a phase during which the universe’s dynamics are dominated by the kinetic energy of a rolling scalar field, which may be the inflaton or another scalar. This phase, referred to as kination, was first proposed in [4], characterised by equation-of-state parameter $w=1$. For a more comprehensive review of kination we refer the reader to [5]. Kination is especially relevant in the context of string theory, where ultra-light moduli fields can dominate early-universe dynamics [6]. String cosmology models involving either the inflaton or other moduli may naturally give rise to kination if the relevant modulus potential features a steep region at the end of inflation–where the potential decreases exponentially, enabling the field to acquire significant kinetic energy–followed by a flat region where the field eventually slows down. What happens afterward depends on the exact shape of the potential.
[…]
@@
The existence of a long kination era within the cosmological timeline may leave imprints in cosmological data that could be searched for in the foreseeable future [20, 21, 22]. However, the authors of [23, 17, 24] pointed out that, according to perturbation theory, such a kination era is expected to be unstable due to the growth of scalar-field perturbations. Unlike the homogeneous background field, these perturbations behave like a relativistic perfect fluid with an equation-of-state parameter of $w=1/3$. The authors of [24] showed that such scalar perturbations may act as a background for the homogeneous scalar field, which ends up tracking its own perturbation and eventually behaves like radiation. Generally speaking, such considerations suggest that a kination era may not exceed 1011$\sim$ 10-11∼ 10 - 11 e-folds  [23, 17]. In [17, 23, 24], the fate of small scalar perturbations is explored in the context of an initially kinetic-energy-dominated background where the gravitational backreaction is ignored. Perturbation theory has provided valuable insights into this regime; however, its main limitation is that it cannot capture strong gravitational effects or backreaction.
@@
This paper is organised as follows: In Section II we provide an introductory framework on kination and our numerical simulation setups, in Section III we present our numerical results for sub-horizon and super-horizon perturbations respectively, focusing on the conditions for black hole formation in the super-horizon case. In Section IV we discuss cosmological implications for the production of primordial black holes in cosmology, and we conclude in Section V. In this work, we use the non-reduced Planck mass, $m_{P}^{2}\equiv\hbar c/G=G^{-1}$ with $\hbar=c=1$ for theoretical analysis, and we use geometric units with $G=c=1$ for simulation works.
II Theory and Methodology
II.1 Perturbations in Kination Epoch
@@
$S=\int d^{4}x\sqrt{-g}\left(\frac{m_{P}^{2}}{16\pi}R-\frac{1}{2}(\nabla\phi)^{2}-V(\phi)\right)\,,\hskip-14.22636pt$ (1)
@@
$\ddot{\phi}+3H\dot{\phi}-\nabla^{2}\phi=0\,.$ (2)
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi}{3m_{P}^{2}}\left(\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}(\nabla\phi)^{2}\right)\,.$ (3)
Before fully solving Einstein’s equations and investigating the effect of non-linear dynamics on the system, let us briefly recap basic results from cosmological perturbation theory, in which the full field solution $\phi(t,x)\equiv\bar{\phi}(t)+\delta\phi(t,\vec{x})$ is decomposed into a homogeneous background solution and small spatial perturbations. The background energy density consists purely of kinetic energy, such that
$\overline{\rho}\equiv\frac{1}{2}\dot{\bar{\phi}}^{2}\propto a^{-6}~.$ (4)
Meanwhile, the linear perturbations behave as $\rho_{\delta\phi}\propto a^{-4}$ in the sub-horizon regime, $k\gg H$, and $\rho_{\delta\phi}\propto a^{-2}$ in the super-horizon regime, $k\ll H$. For more details on perturbation theory, we refer the reader to Appendix A.
In the homogeneous limit where $\delta\phi=0$, $\rho\equiv\overline{\rho}\propto a^{-6}$, until another fluid with a different equation-of-state parameter starts dominating. However, in the presence of perturbations, the situation becomes more interesting because the energy density of the background field redshifts faster than the energy density of its own perturbations. Whenever a notion of a homogeneous background field can be defined—as is the case in the presence of Gaussian and deeply sub-horizon perturbations– the metric evolution is still well described by standard FLRW cosmology. This could lead to the domination of the perturbations, and possibly to tracker solutions in the presence of a subdominant potential, as shown in [24].
@@
Depending on the abundance and amplitude of super-horizon perturbations, the formation of such primordial black holes (PBHs) may play an important role in cosmology, as it could lead to a period of early matter domination after kination with observational consequences. The critical value of the overdensity associated with a curvature perturbation at the time it enters the horizon, $\delta_{c}$, is essential to the calculation of such a PBH abundance (see e.g. [25, 26]). The value of this critical overdensity was studied analytically using a three-zone solution (a spherically symmetric top-hat overdensity surrounded by a matching region, itself embedded in an FLRW universe) by [27] and has the following form
$\delta_{c}=\frac{3(1+w)}{5+3w}\sin^{2}\left(\frac{\pi\sqrt{w}}{1+3w}\right)\,,$ (5)
@@
II.2 Initial Scalar Data
Our simulation domain is a 3-dimensional box with length $L$ and periodic boundary conditions. We adopt a “pseudo-isotropic” setup, with a sinusoidal inhomogeneity of the scalar field in all three spatial directions, while keeping the kinetic background homogeneous:
$\phi_{0}=\bar{\phi}_{0}+\frac{\Delta\phi}{3}\sum^{3}_{i=1}\cos(\frac{2\pi N_{\phi,0}x_{i}}{L})\,,$ (6)
$\dot{\phi}_{0}=\sqrt{\frac{3}{4\pi}}m_{P}H_{0}$ (7)
where the subscript “0” denotes a quantity evaluated at initial time, $x_{i}$ labels the spatial coordinates on the hyperslice evolving by the time coordinate $t$, $\bar{\phi}_{0}$ is the initial homogeneous scalar field, $\Delta\phi$ is the amplitude of the sinusoidal inhomogeneities, $N_{\phi,0}$ is the mode number of the initial scalar field inhomogeneity, $H_{0}$ is the initial Hubble parameter in the absence of the inhomogeneity, i.e. in pure kination limit. In this work, we have chosen to study the fate of a single self-perturbation mode of a scalar field. A thorough treatment of a full spectrum of modes and the effects of kinetic inhomogeneity is beyond the scope of this paper and is left for future investigation.
[…]
$L\equiv(N_{\phi,0}H_{0})^{-1}=(N_{\phi,0}\dot{\phi}_{0})^{-1}\sqrt{\frac{3}{4\pi}}m_{P}\,.$ (8)
With this setup, the gradient energy density, defined as
$\rho_{\nabla}\equiv\frac{1}{2}\gamma^{ij}\partial_{i}\phi\partial_{j}\phi\,.$ (9)
will be the sole contributor to the initial perturbation energy density. We can compute its average at initial time to be
$\langle\rho_{\nabla,0}\rangle=\frac{\Delta\phi^{2}N_{\phi,0}^{2}\pi^{2}}{3L^{2}}\,,$ (10)
[…]
$\langle X\rangle\equiv\frac{\int_{\Sigma}d^{3}x\sqrt{\gamma}X}{\int_{\Sigma}d^{3}x\sqrt{\gamma}}\,.$ (11)
We introduce the non-linearity parameter
$b\equiv\frac{\langle\rho_{\nabla,0}\rangle}{\overline{\rho}_{0}}\,,$ (12)
with
$\overline{\rho}_{0}\equiv\frac{1}{2}{\dot{\phi}}_{0}^{2}$ (13)
which indicates the non-linearity of the initial inhomogeneity and allows us to tune the amplitude of the scalar field inhomogeneity accordingly.
II.3 Initial Geometric Data
@@
$ds^{2}=-\alpha^{2}dt^{2}+\gamma_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt)\,,$ (14)
@@ II.4 Evolution and Diagnostics
$\partial_{t}\alpha=-2\alpha(K-\langle K\rangle-2\Theta)+\beta^{i}\partial_{i}\alpha,$ (17)
This gauge choice maintains the lapse function roughly at its initial value for relatively homogeneous expanding regions while approaching the 1+1+1 +log form in collapsing regions with $\langle K\rangle<0$. It thus better regularises the behaviour of the lapse function in a cosmological spacetime. We adopt the standard Gamma-driver condition
$\partial_{t}\beta^{i}=\frac{3}{4}B^{i},$ (18)
$\partial_{t}B^{i}=\partial_{t}\tilde{\Gamma}^{i}-B^{i}.$ (19)
In this work, we focus on studying the evolution of the scalar field in the presence of a self-perturbation. Inhomogeneities in the form of gradient energy induce gravitational backreaction and result in kinetic inhomogeneities. This also contribute to the perturbation energy density. To study how the total perturbation energy density behaves, we need to identify both the gradient and kinetic perturbation energy densities.
@@
$\Pi\equiv\frac{1}{\alpha}(\dot{\phi}-\beta^{i}\partial_{i}\phi_{0}).$ (20)
@@
$\langle\rho_{\delta\phi}\rangle\equiv\langle\rho_{\nabla}\rangle+\langle\rho_{K}\rangle-\langle\overline{\rho}\rangle,$ (21)
@@
$\langle\rho_{K}\rangle\equiv\frac{1}{2}\langle\Pi^{2}\rangle,$ (22)
$\langle\overline{\rho}\rangle\equiv\frac{1}{2}\langle\Pi\rangle^{2}.$ (23)
This practically means the perturbation energy density is equal to the average gradient density plus the variance of the kinetic energy.
@@
We calculate the masses of the black holes by measuring the minimum area of the apparent horizon. As there is no initial angular momentum in the system, we can approximate the black hole mass with the Schwarzschild radius, $M_{\text{BH}}^{2}\approx r^{2}_{\text{BH}}/4\approx A_{\text{AH}}/16\pi$. In geometric units, $[M]\sim[L]\sim[T]$. We will denote the Hubble scale at the start of the kination period as $H_{\text{kin},0}$. We define our geometric mass according to [36]
$H_{\text{kin},0}=H_{0}\left(\frac{Mg}{m_{P}}\right)^{n},$ (24)
@@
with $H_{0}=1/(N_{\phi,0}L)$ defined in Eq. (8), and $n=-1$ denotes the dimensionality of $H$ in geometric units. The masses of the black holes in Planck units are then converted in a similar fashion with
$M_{\text{BH,pl}}=M_{\text{BH,g}}\left(\frac{M_{g}}{m_{P}}\right).$ (25)
III Numerical Results
@@
For our simulations, we use the multi-purpose numerical relativity code GRChombo [37, 31]. We tune the initial inhomogeneity amplitude and the field momentum based on the ratio $b$ defined in Eq. (12). Our goal is to investigate the fate of the spacetimes when large scalar inhomogeneities emerge in a kination background. We present our results based on the initial characteristic wavelengths of the inhomogeneities. Perturbations with $\lambda_{0}<H_{0}^{-1}$ constitute initially sub-horizon inhomogeneities and conversely, $\lambda_{0}>H_{0}^{-1}$, super-horizon inhomogeneities.
III.1 Sub-Horizon and Near-Horizon Scale Perturbations

### v2

@@ 1 Introduction
The standard cosmological narrative begins with a period of slow-roll inflation [30, 35, 2], during which the dynamics are governed by a scalar field—the inflaton—rolling down a nearly flat potential. This phase provides solutions to the horizon and flatness problems, while also stretching quantum fluctuations of the inflaton field, which later seed the formation of large-scale structures. After inflation ends, it is usually assumed that the inflationary sector reheats the universe by transferring most of its energy to Standard Model particles. This marks the beginning of the Hot Big Bang phase, initially characterized by a radiation-dominated plasma with an equation-of-state parameter $w=1/3$, which eventually transitions to matter domination with $w=0$, as the universe cools. Interestingly, there are no strong constraints on the equation of state of the universe during the period between the end of inflation and the onset of reheating. This opens up the possibility for a variety of cosmological scenarios.
One such possibility is that the post-inflationary–but pre-reheating–epoch is a phase during which the universe’s dynamics are dominated by the kinetic energy of a rolling scalar field, which may be the inflaton or another scalar. This phase, referred to as kination, was first proposed in [34], characterised by equation-of-state parameter $w=1$. For a more comprehensive review of kination we refer the reader to [29]. Kination is especially relevant in the context of string theory, where ultra-light moduli fields can dominate early-universe dynamics [15]. String cosmology models involving either the inflaton or other moduli may naturally give rise to kination if the relevant modulus potential features a steep region at the end of inflation–where the potential decreases exponentially, enabling the field to acquire significant kinetic energy–followed by a flat region where the field eventually slows down. What happens afterward depends on the exact shape of the potential.
Suppose the potential is of a runaway form, the field slows down via Hubble friction, possibly initiating a secondary period of slow-roll inflation. It is worth noting that this feature is common in quintessential inflation models [38, 37, 19, 11, 22, 23] and more generally, in non-oscillatory inflation scenarios [25, 33], where the inflaton field undergoes $(i)$ a slow-roll phase in the early universe, $(ii)$ a phase of acceleration as the potential decreases sharply and the universe enters into a kination era, and $(iii)$ a second phase of slow-roll during which the field decelerates to eventually either behave like a cosmological constant or track the universe’s background energy density at late times [43, 41, 25, 33, 5]. On the other hand, if the potential features a local minimum, the field may eventually settle into it, acquiring a finite mass and oscillating before decaying to reheat the universe, as occurs in the so-called Large Volume Scenarios (LVS) [10, 18, 5].
@@
The existence of a long kination era within the cosmological timeline may leave imprints in cosmological data that could be searched for in the foreseeable future [26, 28, 40]. However, the authors of [eröncel2025universalbounddurationkination, 5, 36] pointed out that, according to perturbation theory, such a kination era is expected to be unstable due to the growth of scalar-field perturbations. Unlike the homogeneous background field, these perturbations behave like a relativistic perfect fluid with an equation-of-state parameter of $w=1/3$. The authors of [36] showed that such scalar perturbations may act as a background for the homogeneous scalar field, which ends up tracking its own perturbation and eventually behaves like radiation. Generally speaking, such considerations suggest that a kination era may not exceed $\sim 10-11$ e-folds  [eröncel2025universalbounddurationkination, 5].
In [5, eröncel2025universalbounddurationkination, 36], the fate of small scalar perturbations is explored in the context of an initially kinetic-energy-dominated background where the gravitational backreaction is ignored. Perturbation theory has provided valuable insights into this regime; however, its main limitation is that it cannot capture strong gravitational effects or backreaction.
@@
This paper is organised as follows: In Section 2 we provide an introductory framework on kination and our numerical simulation setups, in Section 3 we present our numerical results for sub-horizon and super-horizon perturbations respectively, focusing on the conditions for black hole formation in the super-horizon case. In Section 4 we discuss cosmological implications for the production of primordial black holes in cosmology, and we conclude in Section 5. In this work, we use the non-reduced Planck mass, $m_{P}^{2}\equiv\hbar c/G=G^{-1}$ with $\hbar=c=1$ for theoretical analysis, and we use geometric units with $G=c=1$ for simulation works.
2 Theory and Methodology
2.1 Perturbations in Kination Epoch
@@
[…]
@@
$\ddot{\phi}+3H\dot{\phi}-\nabla^{2}\phi=0\,.$ (2.2)
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi}{3m_{P}^{2}}\left(\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}(\nabla\phi)^{2}\right)\,.$ (2.3)
Before fully solving Einstein’s equations and investigating the effect of non-linear dynamics on the system, let us briefly recap the results from cosmological perturbation theory, in which the full field solution $\phi(t,x)\equiv\bar{\phi}(t)+\delta\phi(t,\vec{x})$ is decomposed into a homogeneous background solution and small spatial perturbations. The background energy density consists purely of kinetic energy, such that
$\overline{\rho}\equiv\frac{1}{2}\dot{\bar{\phi}}^{2}\sim a^{-6}~.$ (2.4)
Meanwhile, the linear perturbations behave as $\rho_{\delta\phi}\sim a^{-4}$ in the sub-horizon regime, $k\gg H$, and $\rho_{\delta\phi}\sim a^{-2}$ in the super-horizon regime, $k\ll H$. For more details on perturbation theory, we refer the reader to Appendix A.
In the homogeneous limit where $\delta\phi=0$, $\rho\equiv\overline{\rho}\sim a^{-6}$, until another fluid with a different equation-of-state parameter starts dominating. However, in the presence of perturbations, the situation becomes more interesting because the energy density of the background field redshifts faster than the energy density of its own perturbations. Whenever a notion of a homogeneous background field can be defined—as is the case in the presence of Gaussian and deep sub-horizon perturbations– the metric evolution is still well described by standard FLRW cosmology. This could lead to the domination of the perturbations as a new cosmological background radiation fluid, and possibly to tracker solutions in the presence of a subdominant potential, as shown in [36].
@@
Depending on the abundance and amplitude of super-horizon perturbations, the formation of such primordial black holes (PBHs) may play an important role in cosmology, as it could lead to a period of early matter domination after kination with observational consequences. The critical value of the overdensity associated with a curvature perturbation at the time it enters the horizon, $\delta_{c}$, is essential to the calculation of such a PBH abundance (see e.g. [32, 12]). The value of this critical overdensity was studied analytically using a three-zone solution (a spherically symmetric top-hat overdensity surrounded by a matching region, embedded in an FLRW universe) by [31] and has the following form
$\delta_{c}=\frac{3(1+w)}{5+3w}\sin^{2}\left(\frac{\pi\sqrt{w}}{1+3w}\right)\,$, (2.5)
@@
2.2 Initial Scalar Data
Our simulation domain is a 3-dimensional box with a length $L$ and periodic boundary conditions. We adopt a “pseudo-isotropic” setup with a sinusoidal inhomogeneity of the scalar field and its velocity in all three spatial directions while keeping the kinetic background homogeneous:
$\phi_{0}=\bar{\phi}_{0}+\frac{\Delta\phi}{3}\sum^{3}_{i=1}\cos(\frac{2\pi N_{\phi,0}x_{i}}{L})\,$, (2.6)
$\dot{\phi}_{0}=\sqrt{\frac{3}{4\pi}}m_{P}H_{0}+\frac{\Delta\dot{\phi}}{3}\sum^{3}_{i=1}\cos(\frac{2\pi N_{\dot{\phi},0}x_{i}}{L}+\theta)$ (2.7)
where the subscript 0 denotes a quantity evaluated at initial time, $x_{i}$ labels the spatial coordinates on the hyperslice evolving by the time coordinate $t$, $\bar{\phi}_{0}$ is the initial homogeneous scalar field, $\Delta\phi$ and $\Delta\dot{\phi}$ are the amplitudes of the sinusoidal scalar and kinetic inhomogeneities, $N_{\phi,0}$ and $N_{\dot{\phi},0}$ are the mode numbers of the initial scalar field and kinetic inhomogeneity, $H_{0}$ is the initial Hubble parameter calculated in the absence of the inhomogeneity, i.e. in the pure kination limit, $\theta$ is the phase difference between the scalar and kinetic inhomogeneities. In this work, we have chosen to study the fate of a single self-perturbation mode of a scalar field. A thorough treatment of a full spectrum of modes is beyond the scope of this paper and is left for future investigation.
Since $V\ll\frac{1}{2}\dot{\phi}^{2}$ during a kination era, we consider the simple case where $V(\phi)=0$ [^1] Note that for this reason, we would not observe in our work the tracker solution described in [36] due to the absence of a potential.. Therefore, the dynamics are insensitive to translations in field space, and we are allowed to choose the initial background field value to be $\bar{\phi}_{0}=0$. Since we impose periodic boundary conditions, the size of the simulation domain, $L$, is chosen to be the initial scalar field inhomogeneity wavelength, which is related to the initial Hubble length via
$L\equiv(N_{\phi,0}H_{0})^{-1}=(N_{\phi,0}\dot{\phi}_{0})^{-1}\sqrt{\frac{3}{4\pi}}m_{P}\,.$ (2.8)
With this setup, we define the gradient energy density as
$\rho_{\nabla}\equiv\frac{1}{2}\gamma^{ij}\partial_{i}\phi\partial_{j}\phi\,.$ (2.9)
We can compute the volume averages at the initial time to be
$\langle\rho_{\nabla,0}\rangle=\frac{\Delta\phi^{2}\pi^{2}}{3L^{2}}\,$, (2.10)
and
$\langle\rho_{\delta K,0}\rangle\equiv\langle\frac{1}{2}\langle\dot{\phi}_{0}^{2}\rangle-\frac{1}{2}\langle\dot{\phi}_{0}\rangle^{2}\rangle=\frac{1}{12}(\Delta\dot{\phi})^{2}$ (2.11)
where $\langle\rangle$ denotes the volume average over the spatial hyperslice, defined as
$\langle X\rangle\equiv\frac{\int_{\Sigma}d^{3}x\sqrt{\gamma}X}{\int_{\Sigma}d^{3}x\sqrt{\gamma}}\,.$ (2.12)
We define the initial overdensity parameter as
$\delta_{0}\equiv\frac{\langle\rho_{\nabla,0}\rangle+\langle\rho_{\delta K,0}\rangle}{\overline{\rho}_{0}}\,$, (2.13)
which indicates the non-linearity of the initial inhomogeneity and allows us to set the amplitude of the inhomogeneities accordingly. In this work, we only consider the presence of one type of initial perturbation at a time and leave the study of initial perturbation mixing for future works.
2.3 Initial Geometric Data
@@
$ds^{2}=-\alpha^{2}dt^{2}+\gamma_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt)\,$, (2.14)
@@ 2.4 Evolution and Diagnostics
$\partial_{t}\alpha=-2\alpha(K-\langle K\rangle-2\Theta)+\beta^{i}\partial_{i}\alpha,$ (2.17)
This gauge choice maintains the lapse function roughly at its initial value for relatively homogeneous expanding regions while approaching the 1+1+log form in collapsing regions with $\langle K\rangle<0$. It thus better regularises the behaviour of the lapse function in a cosmological spacetime. We adopt the standard Gamma-driver condition
$\partial_{t}\beta^{i}=\frac{3}{4}B^{i},$ (2.18)
$\partial_{t}B^{i}=\partial_{t}\tilde{\Gamma}^{i}-B^{i}.$ (2.19)
@@
$\Pi\equiv\frac{1}{\alpha}(\dot{\phi}-\beta^{i}\partial_{i}\phi).$ (2.20)
@@
$\langle\rho_{\delta\phi}\rangle\equiv\langle\rho_{\nabla}\rangle+\langle\rho_{\delta K}\rangle,$ (2.21)
@@
$\langle\rho_{\delta K}\rangle\equiv\langle\rho_{K}\rangle-\overline{\rho}$ (2.22)
$\langle\rho_{K}\rangle\equiv\frac{1}{2}\langle\Pi^{2}\rangle,$ (2.23)
$\overline{\rho}\equiv\frac{1}{2}\langle\Pi\rangle^{2}.$ (2.24)
@@
We calculate the masses of the black holes by measuring the minimum area of the apparent horizon. As there is no initial angular momentum in the system, we can approximate the black hole mass with the Schwarzschild radius, $M_{\text{BH}}^{2}\approx r^{2}_{\text{BH}}/4\approx A_{\text{AH}}/16\pi$. In geometric units, $M]\sim[L]\sim[T]$. We will denote the Hubble scale at the start of the kination period as $H_{\text{kin},0}$. We define our geometric mass according to [7]
$H_{\text{kin},0}=H_{0}\left(\frac{Mg}{m_{P}}\right)^{n},$ (2.25)
@@
with $H_{0}=1/(N_{\phi,0}L)$ defined in Eq. (2.8), and n=1n=-1 denotes the dimensionality of $H$ in geometric units. The masses of the black holes in Planck units are then converted in a similar fashion with
$M_{\text{BH,pl}}=M_{\text{BH,g}}\left(\frac{M_{g}}{m_{P}}\right).$ (2.26)
3 Numerical Results
@@
For our simulations, we use the multi-purpose numerical relativity code GRChombo [16, 4]. We tune the initial inhomogeneity amplitude and the field momentum based on the $\delta_{c,0}$ defined in Eq. (2.13). Our goal is to investigate the fate of the spacetimes when large scalar inhomogeneities emerge in a kination background. We present our results based on the initial characteristic wavelengths of the inhomogeneities. Perturbations with $\lambda_{0}<H_{0}^{-1}$ constitute initially sub-horizon inhomogeneities and conversely, $\lambda_{0}>H_{0}^{-1}$, super-horizon inhomogeneities.
3.1 Sub-Horizon and Near-Horizon Scale Perturbations